OpenStudy (anonymous):

AB is a line 48 cm in length and C is its midpoint. On AB, AC, CB semicircles are described. Determine the radius of the circle inscribed in the space enclosed by three semi-circles.

OpenStudy (anonymous):

  Now use the sine rule to give (12+r)/sin 30 =24/sin120 ........ r=\[8\sqrt{3}-12\] I think